# Massless higher-spin supermultiplets in 5D harmonic superspace

### v1

@@ 4.2 Unconstrained prepotentials
The prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ and their gauge parameters $\lambda^{Mm_{1}\ldots m_{s-1}}$ obey the constraints (93) and (99), respectively. These constraints may be solved in terms of unconstrained full-superspace and analytic superfields.
@@
As will be shown below, both $\lambda$- and $\rho$-gauge symmetries (111) and (106) play fundamental role in formulation of conserved higher currents.
@@ 4.3 Example: component structure of spin-3 supermultiplet
Component structure of the prepotentials (87) for arbitrary higher spin $s$ is given in Appendix B. In this subsection, for simplicity, we will present the component structure of these prepotentials for the $s=3$ supermultiplet and will show that the constraints and gauge symmetries described above are sufficient to eliminate all non-physical components from these superfields. On the one hand, the study of component field structure in this section follows the same logic as in Sec. 3.6 for the $s=2$ supermultiplet. On the other hand, the analysis presented here demonstrates some complications arising for higher spins which are absent in the $s=2$ case.
Spin-3 supermultiplet is described by the prepotentials $h^{++mn}$, $h^{++5m}$ and $h^{++\hat{\alpha}+m}$ subject to constraints (93) and gauge transformations (95). The analyticity constraints may be explicitly solved in the analytic coordinates (65):
@@
In addition, there are algebraic ($\Gamma$-)tracelessness constraints:
@@ 4.3.1 Bosonic fields
and gauge superfield parameters
@@
Note that $l_{3}^{mn\,p}$ is antisymmetric in the first two indices, $l_{3}^{mn\,p}=l_{3}^{[mn]p}$ while $l_{2}^{mp}$ is a general tensor with no symmetry constraints. Additional relations between these components arise from the constraint $(\Gamma_{m})_{\hat{\alpha}\hat{\beta}}l^{\hat{\beta}\hat{\gamma}m}=0$:
@@
Using the gauge freedom of the tensors $l_{1}^{m}$, $l_{2}^{mn}$ and $l_{3}^{mn\,p}$ we can gauge away remaining non-physical degrees of freedom in the prepotentials:
@@
    Finally, it is possible to combine $B_{1}^{pmn}$ and $B_{2}^{m}$ into a totally symmetric tensor ${\boldsymbol{B}}^{pmn}$ with no tracelessness condition imposed and standard spin-3 field gauge transformation
@@
    This is a Fronsdal’s spin-3 field with 21 d.o.f. However, ${\boldsymbol{B}}^{pmn}$ cannot appear as a component in the superfield potential $h_{A}^{++mn}$, whereas the correct component field containing spin-3 degrees of freedom is given by Eq. (131).
  • The auxiliary component $V^{mn\,ij}$ is symmetric both in Lorentz and $SU(2)_{R}$ indices. It contains 42 d.o.f. The auxiliary field $D^{m\,ij}$ adds up 15 d.o.f.
@@ 4.3.2 Fermionic fields
Here $\psi^{\hat{\alpha}\,mn\,i}$ is symmetric and traceless in the Lorentz indices with 112 independent components while the unconstrained field $\chi^{\hat{\alpha}\,m\,i}$ has 40 components. The auxiliary component $\rho^{\hat{\alpha}\,m\,i}$ contains 32 d.o.f. since it is $\Gamma$-traceless, $(\Gamma_{m})_{\hat{\alpha}\hat{\beta}}\rho^{\hat{\beta}\,m\,i}=0$, as a corollary of Eq. (119). The last two terms in Eq. (135a) may be eliminated by a redefinition of $\psi^{\hat{\alpha}\,mn\,i}$, but we keep them to simplify gauge variation of $\psi^{\hat{\alpha}\,mn\,i}$ below.
@@
This demonstrates that the Ansatz for the prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ introduced in Sec. 4.1 is necessary and sufficient to describe irreducible massless spin-3 supermultiplet. A similar analysis for arbitrary integer $s$ is given in Appendix B.
@@
$\delta\phi^{\hat{\alpha}\,i}=$ $\frac{2}{5}\partial_{m}\epsilon^{\hat{\alpha}\,m\,i}\,.$ (143)
@@ 4.4 Gauge-invariant action
Next, we switch to the analytic superspace measure by the rule (31) and apply the following corollaries of the analyticity properties of prepotentials (93) and gauge parameters (99)
@@
Unlike the spin-2 case considered in the previous section, the expressions (148) and (149) have different structure. The former expression, however, may be brought to the form of the latter upon commuting $\Gamma$-matrices, e.g.,
@@
and, modulo an overall normalization coefficient, we obtain the gauge-invariant quadratic action for integer higher spin-$s$ supermultiplet:
@@
We stress that the only conditions required for gauge invariance of the action (153) are the $\Gamma$-tracelessness of the Grassmann-odd prepotential (91) and gauge parameter (98). Although not required for the gauge invariance, the tracelessness conditions of other prepotentials (98) may be consistently imposed and preserved by gauge symmetry. These conditions single out irreducible spin-$s$ supermultiplet.
@@ 4.5 Equations of motion
Although the general variation of the action (153) has moderately simple form,
@@
it is not suitable for obtaining equations of motion because the prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ obey the constraints (93). In Eqs. (103), these constraints are resolved in terms of an unconstrained prepotential $\Psi^{\hat{\alpha}-m_{1}\ldots m_{s-2}}$ in the full superspace and an analytic prepotential $v^{++5m_{1}\ldots m_{s-2}}$ ($D^{+}_{\hat{\alpha}}v^{++5m_{1}\ldots m_{s-2}}=0$). Therefore, the correct gauge-invariant equations of motions for massless higher-spin supermultiplets in 5D harmonic superspace arise from variations of the action (153) over the unconstrained prepotentials:
@@ 4.6 Hypermultiplet coupling and conserved higher currents
A procedure for construction of cubic hypermultiplet vertices with higher-spin supermultiplets in 4D harmonic superspace was developed in Refs. Buchbinder:2022kzl ; Buchbinder:2022vra . The structure of these vertices is essentially fixed by the higher-spin gauge symmetry. Following similar ideas, we find the cubic interaction vertex of the hypermultiplet with the higher spin prepotentials in the 5D harmonic superspace in the following general form:
@@
    The curly braces stand for anticommutator of differential operators. The operator $\hat{H}^{++}$ is similar to $H^{++}$ defined by Eq. (92), but the derivatives $\partial_{m}$ in the last term in Eq. (159) act on both the prepotential $h^{++Mm_{1}\ldots m_{s-2}}$ and the hypermultiplet $q^{+a}$. Integrating by parts these derivatives, it is possible to represent the action (157) in the form
@@
$\lambda$-gauge transformations of the prepotentials (111) imply the conservation conditions of the currents (169) on free hypermultiplet equation of motion:
@@
The $\rho$-gauge symmetry (106) in its turn leads to the following conservation laws:
@@
These equations hold on free hypermultiplet equation of motion (52).
@@ 5 Summary and outlook
In this paper, we constructed quadratic actions for massless higher-spin supermultiplets in 5D harmonic superspace. Within this framework, the higher-spin supermultiplets are described by prepotentials $h^{++Mm_{1}\ldots m_{s-2}}=(h^{++mm_{1}\ldots m_{s-2}},h^{++\hat{\alpha}+m_{1}\ldots m_{s-2}},h^{++5m_{1}\ldots m_{s-2}})$ subject to constraints (93) and gauge transformations (95). These superfields are symmetric and traceless in Lorentz indices while the Grassmann-odd prepotential $h^{++\hat{\alpha}+m_{1}\ldots m_{s-2}}$ is also $\Gamma$-traceless. The latter constraint appears crucial for constructing gauge-invariant actions for $s>2$. These actions are similar to the ones found in Refs. Buchbinder:2021ite ; Buchbinder:2022kzl ; Buchbinder:2022vra for higher-spin supermultiplets in 4D harmonic superspace. In our formulation, two of the three prepotentials (87) are not analytic while the manifest analyticity of prepotentials was enforced in the works BIZ24 ; Buchbinder:2021ite ; Buchbinder:2022kzl ; Buchbinder:2022vra . The advantage of our formulation is the manifest supersymmetry of all constructed superfield actions.
@@
We showed that the analyticity constraints (93) are resolved via unconstrained full-superspace prepotential $\Psi^{\hat{\alpha}-m_{1}\ldots m_{s-2}}$ and an analytic superfield $v^{++5m_{1}\ldots m_{s-2}}$. Superfield equations of motion arise upon variation of the action over these unconstrained prepotentials.
@@
We constructed the gauge-invariant cubic interaction vertex of $q$-hypermultiplet with higher-spin supermultiplets. In the harmonic superspace framework, this interaction naturally arises from covariantization of the harmonic-superspace derivative $D^{++}$, by analogy with the vector multiplet minimal coupling. This cubic vertex allows us to derive higher conserved currents in the free hypermultiplet model.
@@
In the particular case of $s=2$, our superfield action describes the 5D linearized supergravity model with 48+48 component degrees of freedom off shell. In the component field form, this supergravity model was studied in Ref. Zucker:1999ej . Fully non-linear formulation for this variant of supergravity in the harmonic superspace framework remains an open problem. This supergravity theory originates from the conformal supergravity coupled to vector and non-linear multiplets playing the role of conformal compensators. It would be interesting to develop superfield formulations of other variants of 5D supergravity models corresponding to different choices of conformal compensators. We expect that each of these supergravity models can be generalized to a variant formulation of the higher-spin supermultiplets.
@@
In conclusion, we would like to briefly discuss some open problems. First, we note that massless off-shell higher-spin supermultiplets in 5D harmonic superspace can be used to obtain off-shell massive higher-spin supersymmetric theories in four dimensions by performing dimensional reduction. This was our main motivation for the present work. Second, let us point out that our approach can be extended to study conformal higher-spin supermultiplets. In particular, one can build cubic 5D $q$-hypermultiplet interaction vertices with conformal higher-spin supermultiplets, by analogy with the works in 4D case Kuzenko:2021pqm ; BIZ24 . Another important problem is to extend the present formulation to the AdS5 harmonic superspace which was introduced in Ref. Kuzenko:2007aj . [^12] Analytic measure in AdS4 harmonic superspace was explicitly constructed in the recent work Ivanov:2025jdp . Finally, the approach developed in the present paper provides a direction for constructing higher-spin supermultiplets in 6D harmonic superspace because of its similarity to the 5D case.
@@
Linearized supergravity action (36) may be rewritten in terms of ordinary superfields. The constrained prepotentials in (36) can be expressed in terms of $\Psi^{\hat{\alpha}-}$ and $v^{++5}$ according to the relations (26). Next, the gauge freedom allows us to bring $\Psi^{\hat{\alpha}-}$ to the form (365) thus resulting with the unconstrained harmonic-independent superfield $\psi^{\hat{\alpha}i}(z)$. Instead of dealing with the analytic prepotential $v^{++5}$, it suffices to work with a Mezincescu-type prepotential $V^{ij}(z)$ Mezincescu which originates as a gauge-fixed version of $v^{++5}$,
@@
Plugging the expressions (365) and (175) into the action (36), one may evaluate the $u$-integral to result with an integral over the Minkowski superspace. [^13] In the 4D ${\cal N}=2$, there exist different linearized supergravity actions Butter:2010sc ; Butter:2011ym which correspond to different choices of the second compensating supermultiplet. It is plausible that there exist alternatives off-shell formulations for massless higher-spin supermultiplets. The actions for higher-spin supermultiplets may also be recast in terms of ordinary superfields. This will be discussed elsewhere.
@@ Acknowledgements.
SMK acknowledges the kind hospitality of the INFN, Sezione di Padova, the Department of Physics of the University of Mons, and the Department of Physics of the University of Munich during his visits in June–July 2025. This work is supported in part by the Australian Research Council, project No. DP200101944.

### v2

@@ 4.2 Unconstrained prepotentials
The prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ and their gauge parameters $\lambda^{Mm_{1}\ldots m_{s-1}}$ obey the constraints (98) and (104), respectively. These constraints may be solved in terms of unconstrained full superspace and analytic superfields.
@@
As will be shown below, both $\lambda$- and $\rho$-gauge symmetries (116) and (111) play a fundamental role in the formulation of conserved higher currents.
@@ 4.3 Example: component structure of spin-3 supermultiplet
The component structure of the prepotentials (92) for arbitrary higher spin $s$ is given in Appendix B. In this subsection, for simplicity, we will present the component structure of these prepotentials for the $s=3$ supermultiplet and will show that the constraints and gauge symmetries described above are sufficient to eliminate all non-physical components from these superfields. On the one hand, the study of the component field structure in this section follows the same logic as in Sec. 3.7 for the $s=2$ supermultiplet. On the other hand, the analysis presented here demonstrates some complications arising for higher spins which are absent in the $s=2$ case.
The spin-3 supermultiplet is described by the prepotentials $h^{++mn}$, $h^{++5m}$ and $h^{++\hat{\alpha}+m}$ subject to the constraints (98) and gauge transformations (100). The analyticity constraints may be solved explicitly in the analytic coordinates (70):
@@
In addition, there are the algebraic ($\Gamma$-)tracelessness constraints:
@@ 4.3.1 Bosonic fields
and in the gauge superfield parameters
@@
Note that $l_{3}^{mn\,p}$ is antisymmetric in its first two indices, $l_{3}^{mn\,p}=l_{3}^{[mn]p}$ while $l_{2}^{mp}$ is a general tensor with no symmetry constraints. Additional relations between these components arise from the constraint $(\Gamma_{m})_{\hat{\alpha}\hat{\beta}}l^{\hat{\beta}\hat{\gamma}m}=0$:
@@
Using the gauge freedom of the tensors $l_{1}^{m}$, $l_{2}^{mn}$ and $l_{3}^{mn\,p}$ we can gauge away any remaining non-physical degrees of freedom in the prepotentials:
@@
    Finally, it is possible to combine $B_{1}^{pmn}$ and $B_{2}^{m}$ into a totally symmetric tensor ${\boldsymbol{B}}^{pmn}$ with no tracelessness condition imposed and a standard spin-3 field gauge transformation
@@
    This is a Fronsdal’s spin-3 field with 21 d.o.f. However, ${\boldsymbol{B}}^{pmn}$ cannot appear as a component in the superfield prepotential $h_{A}^{++mn}$, whereas the correct component field containing spin-3 degrees of freedom is given by Eq. (136).
  • The auxiliary component $V^{mn\,ij}$ is symmetric in both Lorentz and $SU(2)_{R}$ indices. It contains 42 d.o.f. The auxiliary field $D^{m\,ij}$ adds up 15 d.o.f.
@@ 4.3.2 Fermionic fields
Here $\psi^{\hat{\alpha}\,mn\,i}$ is symmetric and traceless in the Lorentz indices with 112 independent components while the unconstrained field $\chi^{\hat{\alpha}\,m\,i}$ has 40 components. The auxiliary component field $\rho^{\hat{\alpha}\,m\,i}$ contains 32 d.o.f. since it is $\Gamma$-traceless, $(\Gamma_{m})_{\hat{\alpha}\hat{\beta}}\rho^{\hat{\beta}\,m\,i}=0$, as a corollary of Eq. (124). The last two terms in Eq. (140a) may be eliminated by a redefinition of $\psi^{\hat{\alpha}\,mn\,i}$, but we keep them to simplify the gauge variation of $\psi^{\hat{\alpha}\,mn\,i}$ below.
@@
This demonstrates that the Ansatz for the prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ introduced in Sec. 4.1 is necessary and sufficient to describe an irreducible massless spin-3 supermultiplet. A similar analysis for arbitrary integer $s$ is given in Appendix B.
@@
$\delta\phi^{\hat{\alpha}\,i}=$ $,\frac{2}{5}\partial_{m}\epsilon^{\hat{\alpha}\,m\,i}\,.$ (148)
@@ 4.4 Gauge-invariant action
Next, we switch to the analytic superspace measure by the rule (32) and apply the following corollaries of the analyticity properties of the prepotentials (98) and gauge parameters (104)
@@
Unlike the spin-2 case considered in the previous section, the expressions (153) and (154) have a different structure. The former expression, however, may be brought to the form of the latter upon commuting $\Gamma$-matrices, e.g.,
@@
Thus, modulo an overall normalization coefficient, we obtain the gauge-invariant quadratic action for an integer higher spin-$s$ supermultiplet:
@@
We stress that the only conditions required for gauge invariance of the action (158) are the $\Gamma$-tracelessness of the Grassmann-odd prepotential (96) and gauge parameter (103). Although not required for gauge invariance, the tracelessness conditions of other prepotentials (103) may be consistently imposed and preserved by gauge symmetry. These conditions single out an irreducible spin-$s$ supermultiplet.
@@ 4.5 Equations of motion
Although the general variation of the action (158) has a moderately simple form,
@@
it is not suitable for obtaining equations of motion because the prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ obey the constraints (98). In Eqs. (108), these constraints are resolved in terms of an unconstrained prepotential $\Psi^{\hat{\alpha}-m_{1}\ldots m_{s-2}}$ in the full superspace and an analytic prepotential $v^{++5m_{1}\ldots m_{s-2}}$ ($D^{+}_{\hat{\alpha}}v^{++5m_{1}\ldots m_{s-2}}=0$). Therefore, the correct gauge-invariant equations of motions for massless higher-spin supermultiplets in 5D harmonic superspace arise from variations of the action (158) with respect to the unconstrained prepotentials:
@@ 4.6 Hypermultiplet coupling and conserved higher currents
A procedure for the construction of cubic hypermultiplet vertices with higher-spin supermultiplets in 4D harmonic superspace was developed in Refs. Buchbinder:2022kzl ; Buchbinder:2022vra . The structure of these vertices is essentially fixed by the higher-spin gauge symmetry. Following similar ideas, we find the cubic interaction vertex of the hypermultiplet with the higher spin prepotentials in the 5D harmonic superspace in the following general form:
@@
    The curly braces stand for the anticommutator of differential operators. The operator $\hat{H}^{++}$ is similar to $H^{++}$ defined by Eq. (97), but the derivatives $\partial_{m}$ in the last term in Eq. (164) act on both the prepotential $h^{++Mm_{1}\ldots m_{s-2}}$ and the hypermultiplet $q^{+a}$. Integrating by parts these derivatives, it is possible to represent the action (162) in the form
@@
The $\lambda$-gauge transformations of the prepotentials (116) imply the conservation conditions of the currents (174) on the free hypermultiplet equation of motion:
@@
The $\rho$-gauge symmetry (111) leads to the following conservation laws:
@@
These equations hold on the free hypermultiplet equation of motion (57).
@@ 5 Summary and outlook
In this paper, we constructed quadratic actions for massless higher-spin supermultiplets in 5D harmonic superspace. Within this framework, the higher-spin supermultiplets are described by the prepotentials $h^{++Mm_{1}\ldots m_{s-2}}=(h^{++mm_{1}\ldots m_{s-2}},h^{++\hat{\alpha}+m_{1}\ldots m_{s-2}},h^{++5m_{1}\ldots m_{s-2}})$ subject to the constraints (98) and gauge transformations (100). These superfields are symmetric and traceless in Lorentz indices while the Grassmann-odd prepotential $h^{++\hat{\alpha}+m_{1}\ldots m_{s-2}}$ is also $\Gamma$-traceless. The latter constraint appears crucial for constructing gauge-invariant actions for $s>2$. These actions are similar to the ones found in Refs. Buchbinder:2021ite ; Buchbinder:2022kzl ; Buchbinder:2022vra for higher-spin supermultiplets in 4D harmonic superspace. In our formulation, two of the three prepotentials (92) are not analytic while the manifest analyticity of the prepotentials was enforced in the works BIZ24 ; Buchbinder:2021ite ; Buchbinder:2022kzl ; Buchbinder:2022vra . The advantage of our formulation is the manifest supersymmetry of all constructed superfield actions.
@@
We showed that the analyticity constraints (98) are resolved via an unconstrained full superspace prepotential $\Psi^{\hat{\alpha}-m_{1}\ldots m_{s-2}}$ and an analytic superfield $v^{++5m_{1}\ldots m_{s-2}}$. Superfield equations of motion arise upon variation of the action with respect to these unconstrained prepotentials.
@@
We constructed the gauge-invariant cubic interaction vertex of $q$-hypermultiplet with higher-spin supermultiplets. In the harmonic superspace framework, this interaction naturally arises from covariantization of the harmonic superspace derivative $D^{++}$, by analogy with the vector multiplet minimal coupling. This cubic vertex allows us to derive higher conserved currents in the free hypermultiplet model.
@@
In the particular case of $s=2$, our superfield action describes the 5D linearized supergravity model with 48+48 component degrees of freedom off shell. In the component field form, this supergravity model was studied in Ref. Zucker:1999ej . A non-linear formulation for this variant of supergravity in the harmonic superspace framework remains an open problem. This supergravity theory originates from the conformal supergravity coupled to vector and non-linear multiplets playing the role of conformal compensators. It would be interesting to develop superfield formulations of other variants of 5D supergravity models corresponding to different choices of conformal compensators. We expect that each of these supergravity models can be generalized to a variant formulation of the higher-spin supermultiplets.
@@
In conclusion, we would like to briefly discuss some open problems. First, we note that massless off-shell higher-spin supermultiplets in 5D harmonic superspace can be used to obtain off-shell massive higher-spin supersymmetric theories in four dimensions by performing dimensional reduction. This was our main motivation for the present work. Second, let us point out that our approach can be extended to study conformal higher-spin supermultiplets. In particular, one can build cubic 5D $q$-hypermultiplet interaction vertices with conformal higher-spin supermultiplets, by analogy with the works in the 4D case Kuzenko:2021pqm ; BIZ24 . Another important problem is to extend the present formulation to the AdS5 harmonic superspace which was introduced in Ref. Kuzenko:2007aj . [^15] An analytic superspace measure in AdS4 harmonic superspace was explicitly constructed in the recent work Ivanov:2025jdp . Finally, the approach developed in the present paper provides a direction for constructing higher-spin supermultiplets in 6D harmonic superspace because of its similarity to the 5D case.
@@
The linearized supergravity action (37) may be rewritten in terms of ordinary superfields. The constrained prepotentials in (37) can be expressed in terms of $\Psi^{\hat{\alpha}-}$ and $v^{++5}$ according to the relations (27). Next, the gauge freedom allows us to bring $\Psi^{\hat{\alpha}-}$ to the form (381), thus resulting in the unconstrained harmonic-independent superfield $\psi^{\hat{\alpha}i}(z)$. Instead of dealing with the analytic prepotential $v^{++5}$, it suffices to work with a Mezincescu-type prepotential $V^{ij}(z)$ Mezincescu which originates as a gauge-fixed version of $v^{++5}$,
@@
Plugging the expressions (381) and (180) into the action (37), one may evaluate the $u$-integral to obtain an integral over the Minkowski superspace. [^16] In the 4D ${\cal N}=2$ case, there exist different linearized supergravity actions Butter:2010sc ; Butter:2011ym which correspond to different choices of the second compensating supermultiplet. It is plausible that there exist alternative off-shell formulations for massless higher-spin supermultiplets. The actions for higher-spin supermultiplets may also be recast in terms of ordinary superfields. This will be discussed elsewhere.
@@ Acknowledgements.
SMK acknowledges the kind hospitality of the INFN, Sezione di Padova, the Department of Physics of the University of Mons, and the Department of Physics of the University of Munich during his visits in June–July 2025. We are grateful to Jake Stirling for comments on the manuscript. This work is supported in part by the Australian Research Council, project No. DP230101629.